OpenStudy (anonymous):

A farmer has 1235 trees to be planted on a rectangular parcel of land. If there are 24 trees planted in each row and each row must be complete before it is planted, how many trees will be left over after planting?

OpenStudy (anonymous):

51*24 = 1224 ----- # left over = 1235-1224 = 11

OpenStudy (anonymous):

hows you come up with 51

OpenStudy (anonymous):

1235 / 24 = 51.458 divide 1235 by 24 an get 51 but take of .458